# Stretching Beyond the Obvious: A Gradient-Free Framework to Unveil the Hidden Landscape of Visual Invariance

### v1

@@ 3.4 Interpretability of the invariant images by humans and observer networks
To assess whether the invariant images produced by SnS retain sufficient information for correct classification by other visual systems, we tested the ability of humans or other observer neural networks to classify them, using a $12$ alternative forced choice task (AFC). For observer networks, we used several distinct ImageNet-pretrained architectures, both standard (AlexNet, VGG16, ResNet18, ResNet101, ConvNext base, RegNetX-16GF), obtained from PyTorch [34], and robust (ResNet-18 [37], ResNet-50 [37], ConvNext base [38], all $L_{\infty},\ \epsilon=\frac{4}{255}$ ), obtained from RobustBench [39].
We chose $12$ ImageNet categories (goldfish, chickadee, frilled lizard, Dungeness crab, Staffordshire bull terrier, cicada, candle, grand piano, minibus, reflex camera, soccer ball, cup). For each category, we selected 10 reference images $\bm{x}_{\mathrm{ref}}$ and, for each of them, we generated one invariant image by stretching its representation in low_level, mid_level, and high_level in both standard and robust ResNet-50. This yielded $6$ kinds of invariant images per reference, where invariance was always defined relative to the readout unit corresponding to the reference image’s category. For each category, we also included in the classification pool the 10 reference images and 20 MEIs (generated with XDREAM) of the readout unit corresponding to the chosen category (10 for the standard and 10 for the robust network). This yielded a total of 12 categories $\times$ 10 randomly sampled images $\times$ 9 stimulus types (i.e., 1 reference image, 2 MEIs and 6 invariant images). Observer networks had to classify all these $1080$ images; humans had to classify a subset of $540$ images (i.e., $5$ rather than $10$ randomly sampled images per category). Classification accuracy (12-AFC percent correct) was calculated for each of the 9 stimulus types.
To assess human recognizability of invariant images, we recruited 25 human participants using the online Prolific platform. After clicking on a fixation cross at the start of each trial, participants were presented with an image for 50 ms at $\sim$10 of visual angle. After a 20ms blank screen, a backward mask image (random RGB pixel noise) was then presented for 200ms. Images were presented in random order. Participants then clicked one of 12 category buttons that subsequently appeared in a circular arrangement, with a 10-second timeout (Fig.5a). To avoid directional bias, the sequence of category buttons was randomly rotated for each trial. Further details on the task and the statistical analysis are reported in Supplementary Material, Section B.
@@ 4 Results
4.1 SnS: an effective dual formulation for invariance and adversarial attacks
Figure 2: SnS generates effective adversarial and invariant examples. The average activation reduction ($\pm$ SD) of $n=77$ readout units of a robust ResNet50 (relative to their MEIs) is plotted against the $L_{2}$ pixel distance from the MEIs, when the latter were transformed with SnS to yield either adversarial or invariant images, or were subjected to affine transformations (rotation, translation and scaling). The pixel distance refers to $224\times 224$ RGB images with values between 00 and $1$.
@@ 4.1 SnS: an effective dual formulation for invariance and adversarial attacks
SnS generates effective adversarial and invariant images. We first validated the SnS framework by generating invariant and adversarial images for a sample of $77$ readout units of a robust ResNet50 using their MEIs as reference images $\bm{x}_{\star}$, and applying the stretching (to achieve invariance) or the squeezing (to achieve adversarial images) to the pixel representation (see Section 3.1). As shown in Fig. 2, SnS successfully generated effective adversarial examples (top left). These stimuli substantially suppressed the activation of the readout units relative to their MEIs (mean reduction of $111\%\pm 7\%$) being displaced from the MEIs by a mean L2 distance of $72\pm 12$ pixels. Such pixel budget reflects both the network robustness and our stringent unit-silencing objective - a stricter criterion than mere misclassification. Consistent with [40, 18], perturbations were semantically relevant, not noise-like (Fig.2).
@@
The invariant images generated by SnS (bottom right) were also very effective, yielding only a minor drop of the units’ activation relative to the MEIs (mean reduction of $34\%\pm 11\%$). At the same time, they substantially departed from the MEIs, with a mean $L_{2}$ pixel distance ($271\pm 32$ pixels) that considerably exceeded the median distance between ImageNet images, as reported by [18]. More impressively, SnS discovered image transformations that were more extreme (in terms of pixel distance) than those achieved by applying to the MEI standard data augmentations (colored dots/lines), while altering the activation of the readout units substantially less, compared to the most extreme augmentations (darker dots). This indicates that SnS can discover the actual axes of image variation a unit learned to tolerate, exploring invariance manifolds way more precisely than traditional tests with predetermined (e.g., affine) transformations.
@@
Finally, we note that SnS successfully generated invariant images also for units in intermediate convolutional layers (Supplementary Material, Section C). This demonstrates its applicability to the analysis of latent representations that are closer to those of biological neurons.
@@ 4.2 SnS reveals layer-specific invariant manifolds
Figure 3: SnS discovers layer-specific invariances. (a) Example MEI and associated invariant images obtained for a readout neuron (“cup”) in robust ResNet-50 and computed for three different choices of the stretching representational stages (rows). Multiple results are shown for different random initialization seeds (columns) (b) Same as (a) but for a standard ResNet-50. (c) Accuracy of a SVC in discriminating the three classes of invariant images produced by stretching pixel-, mid-, and high-level representations as a function of the number of principal components fed to the classifier.
Figure 4: Representation of the invariant images across the feedforward hierarchy. (a) Normalized average distance between the invariant images generated by SnS and their reference MEIs across the different stages of a robust ResNet50. Different lines indicate the different stretching layers used in the optimization (also reported in the $x$ axis). The average distance between randomly selected natural images from Imagenet (solid gray line) and the average distance between multiple MEIs generated via XDREAM are reported for comparison. (b) Same as (a) for standardly trained network.
@@
Next, we compared the kinds of invariant images that SnS discovered by stretching the representations of the MEIs of the $77$ readout units at different stages of processing along a robust ResNet-50 - i.e., in the pixel space (as already done in Fig. 2), as well as in an intermediate (layer 3, conv 1) and a deep (layer 4, conv 7) convolutional layer. Qualitatively, this yielded radically different sorts of invariances (Fig.3a). Stretching in the pixel-space produced luminance and contrast changes; stretching mid-level representations affected texture and color; stretching high-level representations produced abstract variations like viewpoint changes or multiple object instances. The same qualitative differences were observed when SnS was used to produce invariant images for $77$ readout units of a standard ResNet-50 (Fig.3b), although, in this case, the images appeared to be less interpretable than those obtained for the robust counterpart. Also, the images lacked the high-frequency patterns that are typically found in the MEIs of the units of standard networks [41], suggesting that the SnS generator strongly regularizes the image search towards natural statistics.
@@
To quantify these observations, we applied PCA to the invariant images yielded by SnS. We then used the first $k$ principal components as feature vectors to train an SVC to classify the invariant images according to the layer where the SnS stretching was applied. As shown in Fig.3c, a few components were enough to yield above chance performance (i.e.,   $>33.3\%$ correct), and a few tens of components guaranteed a discrimination accuracy virtually perfect for the standard network (red) and above $80\%$ correct for the robust one (green).
@@
To characterize how the invariant images generated by stretching at a given processing stage were represented across the network, we measured their $L_{2}$ distance from their reference MEIs in the representation provided by each layer of the network (Fig.4; this distance was normalized by the mean, within-category distance of Imagenet images; see Section 3.3). By construction, invariant images that were generated to be maximally different from the MEI in a specific layer had the largest normalized distance in that layer, as compared to invariant images obtained by applying the stretching in other layers. This distance consistently surpassed the one among multiple MEIs of the same unit, demonstrating SnS’s efficacy in exploring larger spans of a unit’s invariance manifold, as compared to searching multiple times for different MEIs. Interestingly, the representational distance between the invariant images and their MEIs was preserved in the layers that were adjacent to the one where the stretching was applied, revealing a clear hierarchical pattern. Stretching in the pixel-space yielded invariant images that remained dissimilar from the MEIs in the first layers of the network; stretching in mid and deep convolutional layers yielded images that were more different from the MEIs in the central and final portion of the network. These hierarchical trends were sharper for the standard network than the robust one, suggesting a smoother, more gradual processing of image transformations.
@@
Hierarchical invariances are robust to representation subsampling. To assess SnS’s applicability to neuroscience, where only a small neuronal subpopulation is typically recorded from a visual area, we generated the invariant images using heavily subsampled hidden layer representations. Both qualitative inspection and representational distance analysis revealed that the resulting invariances closely resembled those obtained using the full layer representation, supporting SnS’s potential for neuroscience experiments (see Supplementary Material, Section D for further details).
@@ 4.3 Invariant images for robust and standard networks are perceptually different
Figure 5: Interpretability of the invariant images by humans and other networks. (a) Illustration of the human classification task with the invariant images generated by SnS for standard and robust networks. (b) Classification accuracies of humans (top; averaged across subjects and categories) and AlexNet (bottom; averaged across categories) displayed for each experimental condition. (c) Classification accuracies across multiple robust (left) and standard (right) networks. (d) Correlation coefficient between average human performances and the performances of each observer model.
We next evaluated how recognizable the invariant images generated by SnS were by humans and other CNNs. Based on [15], we hypothesized that the invariant images for the readout units of a robust network would be more semantically coherent and recognizable than those of a standard network. In our tests, invariant images for both kinds of networks were generated by stretching the low- (pixel), mid- and high-level representations of the same set of natural images sampled from 12 different Imagenet categories (Fig.5a). Human subjects performed a 12 AFC classification task on these invariant images, their references, and MEI controls (Fig.5a, top; see Section 3.4 for a detailed description). Invariant images from the robust network (all generation stages) were significantly more recognizable by humans than their standard network counterparts (Fig.5b, top; p < 0.001, Supplementary Material Table 1 ). Invariant images from the robust network were more recognizable when stretching was applied to earlier-layer representations, with pixel-space stretching yielding the highest discrimination accuracy. Interestingly, an opposite trend was observed for the standard network: invariant images from a late layer were more recognizable than those from a middle layer or from pixel space (p < 0.001 for each of the above comparisons: see Supplementary Material Table 1).
This classification task was replicated with various observer CNNs (both standard and robust). As exemplified by the performances obtained for AlexNet (Fig.5a, bottom), the similarity with the accuracy trends found in humans was remarkable (see Supplementary Material Table 1 ). More in general, we found that the patterns of accuracy of the robust networks were very similar (Fig.5c, left) and highly correlated (Fig.5d) with those of humans. Instead, in the case of the standard networks, smaller models (AlexNet, VGG16, ResNet18) mirrored human behavior, while larger, high-performing models (ResNet101, ConvNeXt-Base, RegNetX-16GF) showed no classification difference between the invariant images obtained for the robust and standard ResNet50 (Fig.5c, right and Fig.5d). Nevertheless, all the networks displayed, as humans did, opposing accuracy trends based on the depth of ResNet50 where the stretching was applied: accuracy decreased across layers for the invariant images generated for robust ResNet50, while it increased in the case of the images generated for the standard ResNet50.
@@ 5 Discussion
We introduced Stretch-and-Squeeze, a novel, model-agnostic framework that reconceptualizes how we probe invariance of visual representations. Instead of testing pre-defined transformations (e.g., affine) or imposing strict representational identity (as with metamers [15]), SnS discovers the actual manifold of transformations a unit tolerates by maximizing stimulus dissimilarity in a chosen representational space while preserving as much as possible the unit’s activation. This exploration of functionally equivalent, yet representationally distinct inputs allows SnS to map a broader, potentially more “ecologically” relevant stretch of a unit’s response manifold and its boundaries. The hierarchical application of SnS shows that CNN units in readout layers are not necessarily that robust to the standard transformations (scaling, translation, etc) often applied to probe invariance [20, 11, 12, 13] (Fig. 2). Rather, they strongly tolerate image changes along radically different axes of abstraction (Fig. 3): from low-level properties (e.g., luminance) to mid-level features (e.g., texture) and, finally, to high-level semantic variations (e.g., object pose).
Our comparison of the invariant images generated for standard and robustly-trained ResNet50 extends prior work on the topic [15] with an innovative characterization. The superior semantic coherence and cross-system recognizability (by humans and other CNNs) of the invariances obtained for robust networks strongly corroborates that adversarial training sculpts representations towards human-aligned perceptual features [41, 18, 15]. More strikingly, standard networks revealed a complex relationship between model capacity and the nature of perceived invariance (Fig.5c). While small standard models often aligned with human judgments, larger, higher-performing standard models demonstrated a surprising ability to recognize idiosyncratic invariances generated from other standard networks, even when these were less understandable to humans or robust models. This is possibly a result of the previously described tradeoff between accuracy and robustness [40]: a model highly optimized for accuracy might learn more brittle invariances, leading to less robust representations.
Another interesting finding, consistent across all human and CNN observers, was the opposite trend for the recognizability of the invariant images generated for the robust and the standard network as a function of stretching depth (Fig.5c). This divergence may stem from fundamental differences in the way these networks process and "digest" image variations, whose representation appears to be diluted across wider stretches of adjacent layers in the robust architecture (see Fig.4).

### v2

@@ 3.2 Interpretability of the invariant images by humans and observer networks
To assess whether the invariant images produced by SnS retain sufficient information for correct classification by other visual systems, we tested the ability of humans or other observer neural networks to classify them, using a $12$-alternative forced choice task (AFC). We used several distinct ImageNet-pretrained architectures as observer networks (ResNet18, VGG16_bn, Wide-ResNet50-2, DenseNet161, ResNeXt50_32x4d, Shufflenet_v2_x1_0), both standard obtained from PyTorch (Paszke et al., 2019), and robust (all $L_{2},\ \epsilon=3$), obtained from Salman et al. (2020). To assess human recognizability of invariant images, we recruited 25 human participants using the online Prolific platform. Details on the procedures to run these interpretability experiments are reported in Supplementary Material, Section B.
@@ 4 Results
4.1 SnS generates effective adversarial and invariant images
@@
We first validated the SnS framework by generating invariant and adversarial images for a sample of $77$ readout units of a $L_{2}$-robust ResNet50 using their MEIs $\bm{x}^{\star}$ as reference images, and applying the stretching (to achieve invariance) or the squeezing (to achieve adversarial images) to the pixel representation (see Section 3.1). For each unit, 10 independent invariant and adversarial images were synthesized, each with a different initialization seed, and distance metrics with respect to the reference MEI were aggregated as the mean over the synthesized images. These metrics were then further aggregated across units as mean $\pm$ SD and reported in Fig. 2. As shown in the figure, SnS successfully generated effective adversarial examples (top left). These stimuli substantially suppressed the activation of the readout units relative to their MEIs (mean reduction of $111\%\pm 7\%$), being displaced from the MEIs by a mean $L_{2}$ distance of $72\pm 12$ pixels. This relatively large pixel budget reflects both the network robustness and our stringent unit-silencing objective - a stricter criterion than mere misclassification. Consistent with Tsipras et al. (2018) and Gaziv et al. (2023), perturbations were semantically relevant, not noise-like (Fig. 2).
@@
The invariant images generated by SnS (bottom right of Fig. 2) were also very effective, yielding only a minor drop of the units’ activation relative to the MEIs (mean reduction of $34\%\pm 11\%$). At the same time, they substantially departed from the MEIs, with a mean $L_{2}$ pixel distance ($271\pm 32$ pixels) that considerably exceeded the median distance between ImageNet images, as reported by Gaziv et al. (2023). More impressively, SnS discovered image transformations that were more extreme (in terms of pixel distance) than those achieved by applying
Figure 2: SnS generates effective adversarial and invariant examples. The average activation reduction ($\pm$ SD) of $n=77$ readout units of a $L_{2}$-robust ResNet50 (relative to their MEIs) is plotted against the $L_{2}$ pixel distance from the MEIs, when the latter were transformed with SnS to yield either adversarial or invariant images, or were subjected to affine transformations (rotation, translation and scaling). The pixel distance refers to $224\times 224$ RGB images with values between 0 and 11.
to the MEI standard data augmentations (colored dots/lines), while altering the activation of the readout units substantially less than the most extreme augmentations (darker dots). This indicates that SnS can discover the actual axes of image variation a unit learned to tolerate, exploring invariance manifolds more precisely than traditional tests with predetermined (e.g., affine) transformations. The same result was obtained when we applied SnS to discover invariant and adversarial images for a subset of readout units, but taking highly effective natural images as references instead of the MEIs (see Fig. S5a). See also Fig. S5b for additional examples of invariant and adversarial images obtained using as references either MEIs or natural images. The ability to obtain invariant and adversarial images is enabled intrinsically by the dual loss formulation of the SnS optimization. Indeed, when images are synthesized using only one of the two loss components at a time, results diverge drastically from those of SnS (see Fig. S6 in Section D of the Supplementary Materials).
@@
Finally, we note that SnS successfully generated invariant images also for units in intermediate convolutional layers (Supplementary Material, Section E). This demonstrates its applicability to the analysis of latent representations that are more analogous to those of biological neurons.
@@ 4.2 SnS reveals layer-specific invariant manifolds
Next, we compared the kinds of invariant images that SnS discovered by stretching the MEI representations of the $77$ readout units at different processing stages along a $L_{2}$-robust ResNet50 - i.e. in the pixel space (as done in Fig. 2), as well as in an intermediate (Layer3_conv1) and a deep (Layer4_conv7) convolutional layer. Qualitatively, this yielded different sorts of invariances (Fig. 3a). Stretching in the pixel-space produced mainly luminance and contrast changes; stretching mid-level representations mainly affected texture and color; stretching high-level representations tended to produce abstract variations like viewpoint changes or multiple object instances. The same qualitative differences were observed when SnS was used to produce invariant images for $77$ readout units of standard ResNet50 (Fig. 3b), although, in this case, the images appeared to be less interpretable than those obtained for the robust counterpart. Also, the images lacked the high-frequency patterns that are typically found in the MEIs of standard networks units (Engstrom et al., 2019b), suggesting that the SnS generator strongly regularizes the image search towards natural statistics.
@@
The qualitative descriptions above are just approximate semantic labels to concisely describe image transformations that are, in fact, too rich and complex to be readily captured by a textual description (e.g., stretching in pixel space also produced size changes). To quantify these observations more objectively, we applied PCA to the invariant images yielded by SnS. We then used the first $k$ principal components as feature vectors to train a Support Vector Classifier (SVC) to classify the invariant images according to the layer where the SnS stretching was applied (for details see Supplementary Material, Section A.4). As shown in Fig. 3c, a few components were enough to yield above chance performance (i.e.,   $33.3\%$ correct), and a few tens of components enabled near-perfect discrimination for the standard network (red) and above $80\%$ correct for the robust network (green).
@@
To characterize how the invariant images generated by stretching at a given processing stage were represented across the network, we measured their $L_{2}$ distance from their reference MEIs in the representation provided by each layer of the network (see Supplementary Material, Section A.5.2 for detail). In addition to demonstrating the effectiveness of SnS to explore invariances far away from their reference MEI and heterogeneous across multiple runs (see Section A.5.2 for detail), the representational distance between the invariant images and their MEIs was preserved in the layers that were adjacent to the one where the stretching was applied, revealing a clear hierarchical pattern (Fig. S3 a,b). Stretching in the pixel-space yielded invariant images that remained dissimilar from the MEIs in the first layers of the network; stretching in mid and deep convolutional layers yielded images that were more different from the MEIs in the central and final portions of the network.
Yet another way to characterize the invariance manifolds is to estimate their intrinsic dimension (ID). We did so for an example readout unit of $L_{2}$-robust ResNet50, using both PCA and 2NN-ID, a state-of-the-art nonlinear estimator (Facco et al., 2017) (see Section H in the Supplementary Material). 2NN-ID yielded ID estimates one order of magnitude lower than PCA, indicating that the geometry of the three manifolds was highly nonlinear (Fig. S10). Both methods, however, revealed a consistent ID ranking across stretching layers: lowest for the pixel space, highest for mid layers, and lower again for deep layers - a pattern mirroring previously reported trends in the ID of image representations across deep CNNs and visual cortex (Ansuini et al., 2019; Muratore et al., 2022a)
Figure 3: SnS discovers layer-specific invariances. (a) Example MEI and associated invariant images obtained for a readout neuron (“cup”) in $L_{2}$-robust ResNet50 and computed for three different choices of the stretching representational stages (rows). Multiple results are shown for different random initialization seeds (columns) (b) Same as (a) but for a standard ResNet50. (c) Accuracy of a SVC in discriminating the three classes of invariant images produced by stretching pixel-, mid-, and high-level representations as a function of the number of principal components fed to the classifier.
@@
Hierarchical invariances are robust to representation subsampling. To assess SnS’s applicability to neuroscience, where only a small neuronal subpopulation is typically recorded from a visual area, we generated the invariant images using heavily subsampled hidden layer representations. Both qualitative inspection and representational distance analysis revealed that the resulting invariances closely resembled those obtained using the full layer representation, supporting SnS’s potential for neuroscience experiments (see Supplementary Material, Section F for further details).
@@ 4.3 Invariant images for robust and standard networks are perceptually different
Figure 4: Interpretability of the invariant images by humans and other networks. (a) Illustration of the human classification task with the invariant images generated by SnS for standard and $L_{2}$-robust networks. (b) Classification accuracies of humans (averaged across subjects and categories) displayed for each experimental condition. (c) Classification accuracies across multiple $L_{2}$-robust (left) and standard (right) networks. (d) Correlation coefficient between average human performances and the performances of each observer model. For each architecture, the darker color indicates the robust model and the lighter color indicates the standard version. (e) Analysis from (c) extended to multiple architectures: ResNet50, ResNet18, and VGG16_bn. Solid lines indicate the average between different observer networks (translucent lines).
We next evaluated how recognizable the invariant images generated by SnS were by humans and other CNNs. In our tests, invariant images for both kinds of networks (i.e., standard and $L_{2}$-robust) were generated by stretching the low- (pixel), mid- and high-level representations of the same set of natural images sampled from 12 different Imagenet categories (Fig. 4a). Human subjects performed a 12-AFC classification task on these invariant images, their reference images, and MEI controls (Fig. 4a, top; for details see Section 3.2 and Supplementary Section B). Invariant images from the robust network (all generation stages) were significantly more recognizable by humans than their standard network counterparts (Fig. 4b, compare the green to the red boxes; p < 0.001, Supplementary Material Table  S2). Invariant images from the $L_{2}$-robust network were more recognizable when stretching was applied to earlier-layer representations, with pixel-space stretching yielding the highest discrimination accuracy. Interestingly, an opposite trend was observed for the standard network: invariant images from a late layer were more recognizable than those from a middle layer or from pixel space (p < 0.001 for each of the above comparisons: see Supplementary Material Table S2). These findings provide a more granular understanding of the perceptual misalignment between humans and CNNs, extending beyond what classic feature visualization tools can reveal with a direct comparison of MEIs. In fact, human participants classified the MEIs obtained for the robust network (Fig. 4b, cyan box) with an accuracy comparable to that achieved on the natural images encoded by the corresponding readout units (Fig. 4b, blue box). By comparison, human classification of the MEIs obtained for the standard network was relatively poor (Fig. 4b, pink box). Therefore, considering only MEIs’ interpretability, the $L_{2}$ robustification process would seem very effective in increasing the alignment of image classification between ResNet50 and human observers. SnS reveals instead that this alignment gradually deteriorates when considering invariant images obtained by stretching the representation in progressively deeper layers, with the interpretability gap between the robust and standard networks progressively narrowing. Interestingly, this also shows how SnS can uncover trends that are complementary to those found by previous approaches aimed at assessing invariance, such as computation of metamers (Feather et al., 2023). Similarly, SnS provides a complementary tool to test model alignment with human perception with respect to previous approaches like controversial stimuli (Golan et al., 2020) and Maximum differentiation competitions (Wang and Simoncelli, 2008), which study alignment through model disagreement. The same applies to other approaches that have examined human interpretability of parametric image transformations at the pixel or patch level, such as Extreme Image Transforms (EITs) (Malik et al., 2023) and Minimal Recognizable Configurations (MIRCs) (Ullman et al., 2016). An interesting question would be to assess how strongly the interpretability of SnS-generated invariant images correlates with the interpretability of EITs across a pool of generator networks, spanning different architectures and training regimes.
The same classification task was administered to various observer CNNs (both standard and $L_{2}$-robust). The alignment with the accuracy trends found in humans was relatively high for all robust observer networks (compare Fig. 4b to 4c,left), while their standard counterparts displayed a more heterogeneous degree of consistency with human choices (compare Fig. 4b to 4c,right). As a result, correlations with human perception were very high and statistically significant for all robust networks, but only for a fraction of the standard ones (Fig. 4d). Similarly to humans, all $L_{2}$-robust observer networks better classified the invariant images obtained for $L_{2}$-robust ResNet50 than for its standard counterpart, with this difference narrowing for higher-order invariances. This demonstrates that, despite the $L_{2}$ robustification process, the invariance landscape of ResNet50 readout units remained idiosyncratic and not fully generalizable to other human and artificial observers. This conclusion held true also when standard networks were used as observers, although, for some models, the interpretability gap between invariant images obtained for robust and standard networks was narrower (Fig. 4c, right). Similar trends were found when SnS was used to uncover the hierarchical invariances of two different subject networks (ResNet18 and VGG16_bn, again $L_{2}$-robust and standard; Fig. 4e).
We also applied SnS to produce hierarchical invariant images for the readout units of $L_{\infty}$-robust ResNet50. In this case as well, we obtained layer-specific invariant manifolds that were easily separable in pixel space (Fig. S11a and b). However, unlike $L_{2}$-robustification, $L_{\infty}$-robustification yielded invariant images whose interpretability by $L_{\infty}$-robust observer networks remained consistently high or even increased with stretching depth (Fig. S11c). This indicates that $L_{\infty}$ adversarial training yields invariance manifolds that are, overall, more interpretable than those produced by $L_{2}$ robustification.
To demonstrate the flexibility of the SnS framework, we performed the same interpretability experiment on invariant images evolved for readout units of a vision transformer (ViT). Similar to CNNs, the nature and interpretability of ViT invariances depended on the hierarchical level at which the representation was stretched. Invariances generated from pixel space were substantially less interpretable than those from mid- and high-level representations, which were remarkably similar and more interpretable. This observation aligns with the view that ViTs learn less strictly hierarchical and more globally-integrated features than CNNs (Dosovitskiy et al. (2020); see Fig. S9 in Section G of the Supplementary Material). Taken together, the results of Fig. 4 and Fig. S9 indicate that the level of transferability of the invariances obtained for the readout units of a generator network depends on three key factors: 1) the architecture of the generator (i.e., CNNs vs. vision transformers); 2) its training procedure (robustified vs. standard, in the case of CNNs); and 3) the hierarchical stage at which the representation is stretched to produce the invariant images.
@@ 5 Discussion
We introduced Stretch-and-Squeeze, a novel, model-agnostic framework that reconceptualizes how we probe invariance of visual representations. Instead of testing pre-defined transformations (e.g., affine) or imposing strict representational identity (as with metamers (Feather et al., 2023)), SnS discovers the actual manifold of transformations a unit tolerates by maximizing stimulus dissimilarity in a chosen representational space while preserving the target unit’s activation. This exploration of functionally equivalent yet representationally distinct inputs allows SnS to map a broader, potentially more “ecologically” relevant stretch of a unit’s response manifold and its boundaries. Compared to other approaches aimed at uncovering invariant manifolds (Cadena et al., 2018; Baroni et al., 2023), SnS is distinguished by its gradient-free and model-agnostic formulation, a critical advantage for applications in neuroscience, especially in cases where reliable digital twin models are unavailable. Moreover, differently from these previous approaches, which have been applied to explore only the low-dimensional invariance fields of visual neurons in early processing stages (i.e, primary visual cortex), SnS operates in a highly expressive search space (i.e., the latent code space with dimension $d=4096$), which imposes only a soft constraint on the dimensionality of the invariant manifolds. This allows sufficient expressivity to effectively map the complex invariant fields of high-order units, such as CNNs’ readout neurons, thus characterizing the invariances of the entire model with respect to the object categories it has learned to classify.
The hierarchical application of SnS shows that CNN units in readout layers are not necessarily highly robust to the standard transformations (scaling, translation, etc) often applied to probe invariance (DiCarlo et al., 2012; Goodfellow et al., 2009; Kheradpisheh et al., 2016; Engstrom et al., 2019c) (Fig. 2). Rather, they strongly tolerate image changes along radically different axes of abstraction (Fig. 3): from low-level properties (e.g., luminance) to mid-level features (e.g., texture) and, finally, to high-level semantic variations (e.g., object pose). This illustrates how the invariance is built hierarchically, and how this process is intertwined with the process of building feature selectivity in the first place: while sensitivity to increasingly complex combinations of visual features emerges across CNN layers, from the standpoint of a readout unit, it is the insensitivity to those same, increasingly complex visual properties that is progressively gained. In other words, while a visual system strives to gain selectivity for increasingly complex visual features, it must also become invariant to those same features’ combinations.
Our comparison of the invariant images generated for standard and robustly-trained ResNet50 extends prior work on the topic (Feather et al., 2023) with an innovative characterization. The superior semantic coherence and cross-system recognizability (by humans and other CNNs) of the invariances obtained for robust networks strongly corroborates that adversarial training sculpts representations towards human-aligned perceptual features (Engstrom et al., 2019b; Gaziv et al., 2023; Feather et al., 2023). However, unlike computation of MEIs or metamers (Feather et al., 2023), SnS reveals additional trends that are consistent across all human and CNN observers: there are opposite trends for the interpretability of the invariant images generated for the $L_{2}$-robust and the standard networks as a function of stretching depth (Fig. 4c, e). Specifically, we show that stretching low-level representations generates human-interpretable invariances in robustly trained networks, whereas it produces poorly interpretable invariances in standard networks. Conversely, stretching high-level representations enhances interpretability in standard networks but decreases it in $L_{2}$-robust networks. These trends are different from those obtained by testing the interpretability of metamers (see Table S3). This is not surprising, given that metamers are invariant images that minimize the distance from a target natural image in a given layer, rather than maximizing it, as SnS does. Thus, while SnS pushes the search for invariant images to the boundaries of the invariance manifold, metamers explore a much narrower neighborhood of the target image (see Table S4). Hence the complementary nature of the two approaches.